Determine the freezing point of a solution that co
Answer #1

Number of moles of I2 = Mass/molar mass = 29.5/254 = 0.1161 moles

Molality = Number of moles of solute/mass of solvent in Kg

=> 0.1161/(109 mL * 0.789 g/mL * 10^(-3))

=> 1.35M

Depression in Freezing Point = i * Kf * m

=> 1 * 1.99 * 1.35

=> 2.686C

New freezing point = -114C - 2.686C

=> -116.686C